For each problem, write a system of equations in two or
three variables. Use the method of your choice to solve each
system.
One night the manager of the Sea Breeze Motel rented
5 singles and 12 doubles for a total of $390. The next night
he rented 9 singles and 10 doubles for a total of $412.
What is the rental charge for each type of room?

Found 2 solutions by unlockmath, stanbon:
Answer by unlockmath(1688)

Hello,
Let x represent singles and y be doubles room. From this we can make two equations.
5x+12y=390
9x+10y=412
Now multiply the first equation by -9 and the second equation by 5 to get:
-45x-108y=-3510
45x+50y=2060
Now we can add these to get:
-58y=-1450 Divide each side by -58 to get:
y=25 Plug this answer into the first equation to solve x:
5x+300=390 Subtract 300 from both sides and divide it by 5 to get:
x=18
There we go. $18 for single and $25 for double. Pretty cheap rent, huh?
RJ
A new book might help you out:
www.math-unlock.com

Answer by stanbon(75887)

One night the manager of the Sea Breeze Motel rented
5 singles and 12 doubles for a total of $390. The next night
he rented 9 singles and 10 doubles for a total of $412.
What is the rental charge for each type of room?
-----------------------------------------------------
Equations:
5s + 12d = 390
9s + 10d = 412
---------------------
Multiply 1st by 9 ; Multiply 2nd by 5
45s + 108d = 9*390
45s + 50d = 5*412
----------------------
Subtract the 2nd from the 1st and solve for "d":
58d = 1450
d = $25 (rental rate on doubles)
----
Substitute into 5s + 12d = 390 to solve for "s":
5s + 12(25) = 390
5s = 90
s = $18 (rental rate on singles)
========================================
Cheers,
Stan H.